18. A trolley, while going down an inclined plane, has an acceleration of 2 cm s–2. What will be its velocity 3 s after the start?

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
Initial velocity of trolley, u = 0 cm s–1 Acceleration, a = 2 cm s–2 Time, t = 3 s We know that final velocity, v = u + at = 0 + 2 × 3 cm s–1 Therefore, the velocity of train after 3 seconds = 6 cm s–1